OpenStudy (anonymous):

Find f(x) and g(x) so that the function can be described as y = f(g(x)). y = 8/square root of quantity 2x+4 @agent0smith

OpenStudy (anonymous):

Realize that is equivalent to multiplying 8 by (2x+4)^-1/2

OpenStudy (agent0smith):

\[\Large y = \frac{ 8 }{ \sqrt {2x + 4} }\] You could let g(x) = sqrt(2x+4) and f(x) = 8/x

OpenStudy (anonymous):

16x+32^-1/2

OpenStudy (anonymous):

f(x) = 8, g(x) = square root of quantity two x plus four.

OpenStudy (agent0smith):

f(x) isn't just 8

OpenStudy (anonymous):

8/sqrt x

OpenStudy (agent0smith):

Now you'd have too many square roots. I already posted something that would work above

OpenStudy (anonymous):

eight divided by square root of x., g(x) = 2x + 4

OpenStudy (agent0smith):

That works too.